# An analysis of debris-induced nonlinear phenomena in a positive ion-negative ion (PINI) plasma

### v1

@@ III.1.1 Perturbation with positively charged external debris
These results are entirely in conformity with what we know about formation of debris-induced pinned solitons and DSWs. We now know that in an $e$-$i$ plasma the pinned solitons formed due to moving negatively charged debris is basically a manifestation of the ions trapped in the phase space vortices formed by the IICSI [15], whereas a positively charged debris induces a DSW in the precursor region. So, for $\delta\sim 1$ (vey less concentration of negative ions), the plasma behaves like the usual $e$-$i$ plasma and no pinned solitons can form for a moving positively charged debris. However for $\delta>1$, the negative ions are attracted by the positively charged debris causing a IICSI-like phenomena for negative ions, leading to formation of phase space vortices which in turn cause formation of dark pinned solitons at the debris position.
III.1.2 Comparison with ion-beam dynamics in a PINI plasma
Let us now discuss the case for a positive ion-beam dynamics in a PINI plasma, where the positive ion-beam is produced in situ by accelerating a part of the ion population in the plasma so that overall quasi-neutrality is maintained throughout [3]. This case is interesting as there are several investigations being carried out by different authors to analyze the formation of nonlinear structures in such plasmas [4, 5].
@@ III.1.2 Comparison with ion-beam dynamics in a PINI plasma
Figure 1: Nonlinear wave for a flowing plasma with a positively charged external debris. All panels are drawn in the rest frame of the debris, which is in the middle of a panel, denoted by a vertical dashed line. The shaded vertical regions indicate the extent of the charged debris. The shaded envelopes on the right hand panels indicated the envelopes of the corresponding dark pinned solitons.
which essentially requires that the positive ion beam to be created in situ. As a result, the formation of the positive ion beam requires an overall depletion of the bulk positive ions. The comparative concentration of negative ion is denoted by the quantity $r=n_{-}/n_{T}$, which can be written in terms of $\delta$
@@
where $k=2\pi/\lambda$ with $\lambda$ being the wavelength of the ion-acoustic wave. In the above expression $T_{e}$ is the electron temperature which $\sim 1\,{\rm eV}$ and $m_{+}$ is the mass of the heaviest ion, which is ${\rm Ar}^{+}$, so that $m_{+}\sim 40m_{p}$, where $m_{p}$ is the mass of a proton. Assuming the experimental double-plasma device to be of length $\sim 90\,{\rm cm}$ [3], we can assume the largest possible ion-acoustic wavelength to be $\sim 0.45\,{\rm m}$. With these parameters, we have $\tau_{{\rm IA}}\sim 0.03$ milli second. If we now consider the electron-ion collision time $\tau_{ei}$, we have
@@
where $\ln\Lambda$ is the Coulomb logarithm $\sim 10-12$. For $n_{e}\sim 10^{7}\,{\rm cm}^{-3}$ [3], we have $\tau_{ei}\sim 3.5$ milli second. We now argue that as $\tau_{ei}\gg\tau_{{\rm IA}}$, we can safely assume that in the ion-acoustic timescale, the plasma will not be able to thermalize. As a result a positive ion beam propagating through the plasma, the positive ions of the beam will not be able to thermalize with the bulk plasma and for all practical purpose, the beam will behave as an external debris. It has been observed in N99 that below a certain critical value for beam velocity, increasing $r$ (higher $\delta$) causes the nonlinearity to become negative from positive. However, beyond the critical value of the beam velocity, the nonlinearity remains positive irrespective of the initial pulse (which is reported only at higher $\delta$ in N99). In our case as well, we see a similar behavior as evident from the panels in the left column of Fig.1. We can see that when debris velocity is $\apprle 1.1$, for low $\delta$ $(\sim 1)$, there is a steepening in the leading edge indicating positive nonlinearity while for higher $\delta$, the steepening is on the falling edge, which indicates negative nonlinearity. We note that the ‘positive nonlinearity’ in the case of a compressive or bright soliton is referred with respect to the nonlinear term $A\phi^{(1)}\partial_{\xi}\phi^{(1)}$ [see Eq.(14)] in a usual KdV equation (without the debris term), which is responsible for steepening of the leading edge of a soliton. On the other hand, when the steepening occurs in the falling edge, it is usually termed as ‘negative nonlinearity’. When debris velocity increases (bottom right panel of Fig.1), the negative nonlinearity which was observed for say $\delta=1.5$ and with low $v_{d}$, decreases and get balanced by dispersion giving rise to a bright KdV soliton. These results are also confirmed through FCT simulation, which are being presented in another subsection.
Figure 3: FCT simulation results of propagation of nonlinear wave in a PINI plasma with and embedded external positively charged debris. These results are equivalent to the theoretical cases obtained through fKdV solution shown in Fig.1. The left panel in the above figure corresponds to the left-bottom panel of Fig.1 and the right panel above corresponds to the right-top panel of Fig.1. These results are plotted at the and of $\tau=100\equiv\varepsilon^{3/2}t$.
@@
We should however be cautious not to compare the entire experimental results of N99 with the results obtained in this work, as the external charged debris in this work is a localised charged perturbation with no assumption of quasi-neutrality, while in N99 and other similar experimental works, the beam is an extended structure with definitive wake region. So, we can probably compare only the leading edge dynamics of a beam-plasma interaction with the precursor region of debris-induced nonlinearity.
III.1.3 Comparison with external perturbation in a PINI plasma
@@
Let us now consider a very recent experimental observation of large-amplitude perturbation in a PINI plasma by Pathak and Bailung [16] (P25 thereafter), where a positive pulse is induced externally to a PINI plasma resulting a large-amplitude positive perturbation (up to $70\%$ of equilibrium plasma density), which is then used to study the effect of controlled Landau damping on nonlinear IA wave. This produces a dissipative shock front when dissipation (through Landau damping) becomes large.
@@
Figure 4: FCT simulation results of propagation of nonlinear wave in a PINI plasma with and embedded external negatively charged debris. These results are equivalent to the theoretical cases obtained through fKdV solution shown in Fig.1. The left panel in the above figure corresponds to the left-middle panel of Fig.2 and the right panel above corresponds to the right-top panel of Fig.2. These results are plotted at the and of $\tau=100\equiv\varepsilon^{3/2}t$.
@@
We argue that dynamically, an external perturbation can be considered to be having the same effect as an external charged debris so far as the leading edge (or precursor region) is concerned. So, we can expect a similar behavior of nonlinear wave front for this external perturbation and perturbation induced by an external debris. In this context, we would like to compare our results as shown in the cases for $\delta=1.02$ with $v_{d}=1.0,1.1$ (two bottom panels of the lefthand column in Fig.1) in the precursor regions and the case shown in Fig.4 of P25, which is comparable to our theoretical situation. In both cases, we have a leading shock front with oscillations which are nothing but dispersive shock wave (DSW) when there is no dissipation. Note that Fig.4 of P25 is for the case when Landau damping is negligible with negligible negative ions $(\delta\sim 1.0)$. The situation is however different in the wake region (or trailing region of the pulse). This is due to the fact that experimentally a one-time external perturbation is induced by pushing a signal which then propagates in the form of discontinuity (either as a single soliton as a ramp structure) through the plasma which then transforms into a DSW as it moves. For the case of external debris however, the perturbation remains intact all throughout the life cycle of the nonlinear wave front giving rise to a very definitive signature in the wake region.
III.1.4 Perturbation with negatively charged external debris
@@
We now present the results for a negatively charged debris in a flowing PINI plasma in Fig.2. All the parameters are kept same as in the case for positively charged debris. The shaded regions in all the panels show the extent of the charged debris with the centered dashed line indicating the debris position. As expected, bright pinned solitons are formed for $\delta$ close to unity at $\delta=1.02$, when the relative velocity between debris and PINI plasma $v_{d}$ exceeds a certain value (left-top and middle panels of Fig.2). It can be clearly seen that as $v_{d}$ increases, these pinned solitons finally vanishes even for $\delta\rightarrow 1$. For higher values of $\delta,$ IA waves in the precursor region propagate showing an oscillatory shock excitation which is not fully formed at a lower debris velocity. However, as $v_{d}$ increases amplitude of the precursor wave-front also increases and fully developed DSWs are observed to be formed. Here, it is worth mentioning that, similar behavior was also reported for the propagation of a rarefactive pulse in presence of positive ion beam in high negative ion density plasma [5]. Increasing the beam velocity resulted in a rise in soliton amplitude and a knee formation at the downstream, culminating in the formation of a shock-like structure. A realistic fluid simulation for much longer time period shows that these DSWs, as they propagate away from the site of the debris, get transformed into a train of individual solitons after segregation of the peaks, about which we shall discuss in Sec.IV (Fig.5). The pinned solitons at the site of the debris are deformed due to the presence of the shock. Also, behind the debris, soliton-like structures are seen to be excited due to IA oscillations.
@@ IV Flux-corrected transport (FCT) simulation
To corroborate the above theoretical analysis, we now present an FCT simulation of the PINI plasma with an embedded external charged debris. We use a 1-D multi-fluid FCT (mFCT) code [13] based on Boris’s original algorithm [17] with Zalesak’s flux limiter [18]. The FCT formalism requires the hydrodynamic equations Eq.(1-3) to be put in the form of a generalized continuity equation,
@@
where $f=(n_{+},n_{-},n_{+}v_{+},n_{-}v_{-})$ is the physical quantity to be solved, $(fv)$ is the corresponding flux, and $s$ is the source term. The external charged debris is entered through the Poisson equation Eq.(4) as before which is to be solved at every time step along with Eq.(21). Our FCT simulation results are presented through Fig.3 for the case with positively charged debris and through Fig.4 for negatively charged debris. In Fig.3, we have shown the results of the mFCT simulation for the cases shown in the left-bottom panel of Fig.1 (corresponding to the left panel of Fig.3) and right-top panel of Fig.1 (corresponding to the right panel of Fig.3). These curves are plotted at the end of $\tau=100\equiv\varepsilon^{3/2}t$. The $x$-axis is also respectively scaled for $\xi\equiv\varepsilon^{1/2}x$. Our estimate shows that the effective $\varepsilon\sim 0.7$ for $\rho_{0}=0.5$, which is used in the fKdV solution as well as in the mFCT simulation. In both cases, the external charged debris is a Gaussian profile given by $\rho_{{\rm ext}}(\xi)=\rho_{0}e^{-\xi^{2}/\Delta}$.
@@
We now present the results of FCT simulation for negatively charged debris in Fig.4. All other parameters except the debris charge is same as in the case of Fig.3. The left panel of Fig.4 corresponds to the fKdV solution presented in left-middle panel of Fig.3 and the right panel corresponds to the right-top panel of Fig.3. Though the bright pinned solitons for $\delta=1.02$ is properly captured in the FCT simulation, DSWs for higher $\delta$ is shown in both $v_{d}=1.0$ and $1.5$, while in fKdV solutions, only in the case of $v_{d}=1.5$, these DSWs seem to appear fully formed. Besides, the distinct soliton-like structures in the wake regions present in the fKdV solutions for both cases seem to absent in the simulation. Instead, singly peaked pinned solitons with almost no wake oscillations for $v_{d}=1.0$ and peak-modulated pinned solitons [10] and a turbulent wake trail due to the propagating shock for $v_{d}=1.5$ are found to exist. These discrepancies between the fKdV solution and FCT simulations perhaps can be attributed to the weak nonlinearity approximation that is in-built in fKdV derivation, while in FCT simulation there is no approximation on nonlinearities involved. As a result, nonlinear oscillations and structures tend to be more prominent in the simulation. Nevertheless, we believe that to large extent the fKdV dynamics is similar to the FCT simulations and our observations about similarities between debris-induced nonlinearities and nonlinearities introduced by positive ion-beam and large-amplitude perturbation in a PINI plasma can be fairly ascertained.
@@
It is important to note that the DSWs seen in Fig. 2 and 4 are called ‘quasi-steady state shocks’ [19] as the initial oscillatory shock profile is not retained over a longer time period ($t\geq 200$). Fig.5 shows the potential distribution of one such shock profile at different time-steps $(t=0-100)$. Initially, the external charge perturbation introduces a highly nonlinear IA wave and the leading edge of the precursor becomes steepened which forms a DSW with an oscillatory tail behind due to its dispersive nature. This non-steady shock profile, after a relatively long time $(t\geq 200)$, transforms into a train of individual propagating solitons as nonlinearity gets balanced by dispersion and each and every oscillatory peaks in the DSW separates into individual soliton structures. This effect is already observed and confirmed by earlier investigations on IA shock waves both in theory and experiments [19]. The positive hump observed at the site of the debris is the signature formation of a single-peak bright pinned soliton.

### v3

@@ III.1 Positively charged external debris
$\beta_{e}=\frac{e^{2}}{4\pi\epsilon_{0}RT_{e}},$ (20)
where $e$ is the electronic charge and $R$ is the radius of a spherical debris [32]. This result is rigorously proved for impurity particles in a plasma using the orbit motion limited (OML) theory [33] and numerical simulation, which can be readily interpreted for charged debris as well. This can also be extended to debris with positive charge.
Figure 1: Nonlinear wave for a flowing PINI plasma with a positively charged external debris. All panels are drawn in the rest frame of the debris, which is in the middle of a panel, denoted by a vertical dashed line. The shaded vertical regions indicate the extent of the charged debris. The shaded envelopes on the right hand panels indicated the envelopes of the corresponding dark pinned solitons.
If we now consider the case of charged debris in space plasma in low earth orbit (LEO), the size of the debris can vary from being very small to very large, but anything larger than a few Debye length is capable of inducing nonlinear structures. Considering typical parameters for LEO plasma [34, 35], the value for $\beta_{e}$ comes out to be $\sim 10^{-6}$ even for $R\sim\lambda_{D}$. For ions (positively charged debris), this comes out to be $\beta_{i}\sim 10^{-3}$. In laboratory experiments involving debris $\beta_{e}\sim 10^{-5}$ assuming debris of size $\sim\lambda_{D}$ [36].
These results are entirely in conformity with what we know about formation of debris-induced pinned solitons and DSWs. We now know that in an $e$-$i$ plasma the pinned solitons formed due to moving negatively charged debris is basically a manifestation of the ions trapped in the phase space vortices formed by the IICSI [19], whereas a positively charged debris induces a DSW in the precursor region. So, for $\delta\sim 1$ (vey less concentration of negative ions), the plasma behaves like the usual $e$-$i$ plasma and no pinned solitons can form for a moving positively charged debris.
However for $\delta>1$, the negative ions are attracted by the positively charged debris causing a IICSI-like phenomena for negative ions, leading to the formation of phase space vortices which in turn cause formation of dark pinned solitons at the debris position.
III.1.1 Comparison with ion-beam-induced nonlinearities
Let us now discuss the first case where under certain circumstances, a positive ion-beam in a PINI plasma mimics the responses of a debris-induced nonlinearities. In this experiment, the positive ion-beam is produced in situ by accelerating a part of the ion population in the plasma so that overall quasi-neutrality is maintained throughout [11]. This case is also interesting as there are several investigations being carried out by different authors to analyze the formation of nonlinear structures in such plasmas [12, 13].
Figure 3: FCT simulation results of propagation of nonlinear wave in a PINI plasma with and embedded external positively charged debris. These results are equivalent to the theoretical cases obtained through fKdV solution shown in Fig.1. The left panel (a) in the above figure corresponds to Fig.1(c) and the right panel (b) above corresponds to Fig.1(d). These results are plotted at the and of $\tau=100\equiv\varepsilon^{3/2}t$.
@@ III.1.1 Comparison with ion-beam-induced nonlinearities
which essentially requires that the positive ion beam to be created in situ. As a result, the formation of the positive ion beam causes an overall depletion of the bulk positive ions. The comparative concentration of negative ion is denoted by the quantity $r=n_{-}/n_{T}$, which can be written in terms of $\delta$
@@
where $k=2\pi/\lambda$ with $\lambda$ being the wavelength of the ion-acoustic wave. In the above expression, $T_{e}$ is the electron temperature which is $\sim 1\,{\rm eV}$ and $m_{+}$ is the mass of the heaviest ion, which is ${\rm Ar}^{+}$, so that $m_{+}\sim 40m_{p}$, where $m_{p}$ is the mass of a proton. Assuming the experimental double-plasma device to be of length $\sim 90\,{\rm cm}$ [11], we can assume the largest possible ion-acoustic wavelength to be $\sim 0.45\,{\rm m}$. With these parameters, we have $\tau_{{\rm IA}}\sim 0.03$ milli second. If we now consider the electron-ion collision time $\tau_{ei}$, we have
@@
where $\ln\Lambda$ is the Coulomb logarithm $\sim 10-12$. For $n_{e}\sim 10^{7}\,{\rm cm}^{-3}$ [11], we have $\tau_{ei}\sim 3.5$ milli second. We now argue that as $\tau_{ei}\gg\tau_{{\rm IA}}$, it can be safely assumed that in the ion-acoustic timescale, the plasma will not be able to thermalize. As a result, a positive ion beam propagating through the plasma will not be able to thermalize with the bulk plasma and for all practical purposes, the beam will behave as an external debris. It has been observed in N99 that below a certain critical value for beam velocity, increasing $r$ (higher $\delta$) causes the nonlinearity to become negative from positive. However, beyond the critical value of the beam velocity, the nonlinearity remains positive irrespective of the initial pulse (which is reported only at higher $\delta$ in N99). In our case with debris as well, we see a similar behavior as evident from the panels in the left column of the figure i.e. Fig.1(a), (b), and (c). We can see that when debris velocity is $\apprle 1.1$, for low $\delta$ $(\sim 1)$, there is a steepening in the leading edge indicating positive nonlinearity while for higher $\delta$, the steepening is on the falling edge, which indicates negative nonlinearity. We note that the ‘positive nonlinearity’ in the case of a compressive or bright soliton is referred with respect to the nonlinear term $A\phi^{(1)}\partial_{\xi}\phi^{(1)}$ [see Eq.(14)] in a usual KdV equation (without the debris term), which is responsible for steepening of the leading edge of a soliton. On the other hand, when the steepening occurs in the falling edge, it is usually termed as ‘negative nonlinearity’. When debris velocity increases [Fig.1(f)], the negative nonlinearity which was observed for say $\delta=1.5$ and with low $v_{d}$, decreases and get balanced by dispersion giving rise to a bright KdV soliton. These results are also confirmed through our FCT simulation, which are being presented in Section IV.
@@
We should however be cautious not to compare the entire experimental results of N99 with the results obtained in this work, as the external charged debris in this work is a localized charged perturbation with no assumption of quasi-neutrality, while in N99 and other similar experimental works, the beam is an extended structure with definitive wake region. So, we can probably compare only the leading edge dynamics of a beam-plasma interaction with the precursor region of debris-induced nonlinearity.
III.1.2 Comparison with pulse-induced nonlinearities
@@
Let us now consider a very recent experimental observation of large-amplitude perturbation in a PINI plasma by Pathak and Bailung [24] (P25 thereafter), where a positive pulse is induced externally to a PINI plasma resulting in a large-amplitude positive perturbation (up to $70\%$ of equilibrium plasma density), which is then used to study the effect of controlled Landau damping on nonlinear IAW. This produces a dissipative shock front when dissipation (through Landau damping) becomes large.
@@
Figure 4: FCT simulation results of propagation of nonlinear wave in a PINI plasma with and embedded external negatively charged debris. These results are equivalent to the theoretical cases obtained through fKdV solution shown in Fig.2. The left panel (a) in the above figure corresponds to Fig.2(b) and the right panel (b) above corresponds to Fig.2(d). These results are plotted at the and of $\tau=100\equiv\varepsilon^{3/2}t$.
@@
We argue that dynamically, an external perturbation can be considered to be having the same effect as an external charged debris so far as the leading edge (or precursor region) is concerned. So, we can expect a similar behavior of nonlinear wave front for this external perturbation and perturbation induced by an external debris. In this context, we would like to compare our results as shown in the cases for $\delta=1.02$ with $v_{d}=1.0,1.1$ in Fig.1(b) and (c) in the precursor regions and the case shown in Fig.4 of P25, which is comparable to our theoretical situation. In both cases, we have a leading shock front with oscillations which are nothing but DSWs when there is no dissipation. Note that Fig.4 of P25 is for the case when Landau damping is negligible with negligible negative ions $(\delta\sim 1.0)$. The situation is however different in the wake region (or trailing region of the pulse). This is due to the fact that experimentally a one-time external perturbation is induced by pushing a signal which then propagates in the form of discontinuity (either as a single soliton as a ramp structure) through the plasma which then transforms into a DSW as it moves. For the case of external debris however, the perturbation remains intact all throughout the life cycle of the nonlinear wave front giving rise to a very definitive signature in the wake region.
III.2 Negatively charged external debris
@@
We now present the results for a negatively charged debris in a flowing PINI plasma in Fig.2. All the parameters are kept same as in the case for positively charged debris. The shaded regions in all the panels show the extent of the charged debris with the centered dashed line indicating the debris position. As expected, bright pinned solitons are formed for $\delta$ close to unity at $\delta=1.02$, when the relative velocity between debris and PINI plasma $v_{d}$ exceeds a certain value as shown in Fig.2(a) and (b). It can be clearly seen that as $v_{d}$ increases, these pinned solitons finally vanishes even for $\delta\rightarrow 1$, which is consistent with our observations about pinned solitons in an $e$-$i$ plasma [17, 19]. For higher values of $\delta,$ IAWs in the precursor region propagate showing an oscillatory shock excitation which is not fully formed at a lower debris velocity. However as $v_{d}$ increases, amplitude of the precursor wave-front also increases and fully developed DSWs are observed to be formed. Here, it is worth mentioning that, similar behavior was also reported for the propagation of a rarefactive pulse in presence of positive ion beam in high negative ion density plasma [13]. Increasing the beam velocity resulted in a rise in soliton amplitude and a knee formation at the downstream, culminating in the formation of a shock-like structure. A realistic fluid simulation for much longer time period shows that these DSWs, as they propagate away from the site of the debris, get transformed into a train of individual solitons after segregation of the peaks, about which we shall discuss in Section IV. The pinned solitons at the site of the debris are deformed due to the presence of the shock. Also, behind the debris, soliton-like structures are seen to be excited due to ion-acoustic oscillations.
@@ IV Flux-corrected transport (FCT) simulation
To corroborate the theoretical analysis presented in Section III, we now present an FCT simulation of the PINI plasma with an embedded external charged debris. We use a 1-D multi-fluid FCT code (mFCT) [22] based on Boris’s original algorithm [37] with Zalesak’s flux limiter [38]. The FCT formalism requires the hydrodynamic equations Eq.(1-3) to be put in the form of a generalized continuity equation,
@@
where $f=(n_{+},n_{-},n_{+}v_{+},n_{-}v_{-})$ is the physical quantity to be solved, $(fv)$ is the corresponding flux, and $s$ is the source term. The external charged debris is entered through the Poisson equation Eq.(4) as before which is to be solved at every time step along with Eq.(25). We must however emphasize that the FCT simulation results contain both KdV-type and NLSE-type nonlinear solutions with the latter being less dominant. Nevertheless, in certain parameter regimes, we might see a considerable deviation of the simulation results from fKdV solutions.
@@
Our FCT simulation results are presented through Fig.3 for the case with positively charged debris and through Fig.4 for negatively charged debris. In Fig.3, we have shown the results of the FCT simulation for the fKdV cases shown in Fig.1(c), corresponding to Fig.3(a) and Fig.1(d), corresponding to Fig.3(b). These curves are plotted at the end of $\tau=100\equiv\varepsilon^{3/2}t$. The $x$-axis is also respectively scaled for $\xi\equiv\varepsilon^{1/2}x$. Our estimate shows that the effective $\varepsilon\sim 0.7$ for $\rho_{0}=0.5$, which is used in the fKdV solution as well as in the FCT simulation. In both cases, the external charged debris is a Gaussian profile given by $\rho_{{\rm ext}}(\xi)=\rho_{0}e^{-\xi^{2}/\Delta}$. As we can see that the simulations results are quite comparable to those obtained from the corresponding fKdV solutions, except in the case for $\delta=1.5$ and $v_{d}=1.5$, which can be attributed to slightly different time over which various nonlinearities develop in a simulation and in fKdV solutions.
@@
We now present the simulation results for negatively charged debris in Fig.4. All other parameters except the debris charge is same as in the case of Fig.3. Fig.4(a) corresponds to the fKdV solution presented in Fig.2(b) $(v_{d}=1.0)$ and Fig.4(b) corresponds to Fig.2(d) $(v_{d}=1.5)$. We observe that bright pinned solitons for $\delta=1.02$ and DSWs for higher $\delta$ are properly captured in the FCT simulation, while the DSWs seem to be fully formed in the fKdV solutions only when $v_{d}=1.5$. Besides, the distinct soliton-like structures in the wake regions of the fKdV solutions for both $v_{d}=1.0$ and $1.5$ seem to be absent in the simulation. Instead, singly peaked pinned solitons with almost no wakes for $v_{d}=1.0$ and peak-modulated pinned solitons [17] and a turbulent wake trail due to the propagating shock for $v_{d}=1.5$ are found to exist. These discrepancies between the fKdV solution and FCT simulations perhaps can be attributed to the weak nonlinearity approximation that is in-built in the fKdV derivation, while in FCT simulation there is no approximation on nonlinearities involved. As a result, nonlinear oscillations and structures tend to be more prominent in the simulation, which might also contain modulated wave-packet-like solutions (forced NLSE type). Nevertheless, we believe that to large extent the fKdV dynamics is similar to the FCT simulations and our observations about similarities between debris-induced nonlinearities and nonlinearities introduced by positive ion-beam and large-amplitude perturbation in a PINI plasma can be fairly ascertained.
@@
It is important to note that the DSWs seen in Fig.2 and 4 are called ‘quasi-steady state shocks’ [39] as the initial oscillatory shock profile is not retained over a longer time period ($t\geq 200$). Fig.5 shows the potential distribution of one such shock profile at different time-steps $(t=0-100)$. Initially, the external charge perturbation introduces a highly nonlinear IAW and the leading edge of the precursor becomes steepened which forms a DSW with an oscillatory tail behind due to its dispersive nature. This non-steady shock profile, after a relatively long time $(t\geq 200)$, transforms into a train of individual propagating solitons as nonlinearity gets balanced by dispersion and each and every oscillatory peaks in the DSW separates into individual soliton structures. This effect is already observed and confirmed by earlier investigations on ion-acoustic shock waves both in theory and experiments [39]. The positive hump observed at the site of the debris is the signature formation of a single-peak bright pinned soliton.